There is a famous theorem in Euclidean geometry that states that the sum of
the interior angles of a triangle is 180.
(a) Use the theorem about triangles to determine the sum of the angles of
a convex quadrilateral.
(b) Use the result in Part (1) to determine the sum of the angles of a convex
pentagon.
(c) Use the result in Part (2) to determine the sum of the angles of a convex
hexagon.

Answer by richard1234(7193)

a) Quadrilateral can be divided into two triangles (each with vertices at the vertices of the quadrilateral) so the sum of the angles is 2*180 = 360

b) Same logic as a) but now we can divide into three triangles --> 3*180 = 540

c) 4*180 = 720

We can generalize to n sides and derive the sum of the angles of an n-gon to be 180(n-2).